### v1

# LLM-Enhanced Multimodal Fusion for Cross-Domain Sequential Recommendation

Abstract
Cross-Domain Sequential Recommendation (CDSR) predicts user behavior by leveraging historical interactions across multiple domains, focusing on modeling cross-domain preferences and capturing both intra- and inter-sequence item relationships. We propose LLM-Enhanced Multimodal Fusion for Cross-Domain Sequential Recommendation (LLM-EMF), a novel and advanced approach that enhances textual information with Large Language Models (LLM) knowledge and significantly improves recommendation performance through the fusion of visual and textual data. Using the frozen CLIP model, we generate image and text embeddings, thereby enriching item representations with multimodal data. A multiple attention mechanism jointly learns both single-domain and cross-domain preferences, effectively capturing and understanding complex user interests across diverse domains. Evaluations conducted on four e-commerce datasets demonstrate that LLM-EMF consistently outperforms existing methods in modeling cross-domain user preferences, thereby highlighting the effectiveness of multimodal data integration and its advantages in enhancing sequential recommendation systems. Our source code will be released.
keywords:
Weakly-Supervised Learning, Semantic Segmentation, Large Language Model , Diffusion Model
1 Introduction
Sequential Recommendation (SR) has emerged as a prominent approach for modeling dynamic user preferences by analyzing historical interaction sequences Shani et al. [2002], He et al. [2017a], He & McAuley [2016], Li et al. [2017]. By focusing on predicting the next item of interest based on past behaviors, SR techniques capture the evolving nature of user interests within a single domain. However, SR faces several challenges, including data sparsity, where historical data for many items or users is insufficient, and domain bias, where models are overfitted to a specific domain or item set. These issues limit the ability of SR models to capture the full spectrum of user preferences, particularly when users interact with multiple domains. To overcome these challenges and improve recommendation quality, Cross-Domain Sequential Recommendation (CDSR) Ma et al. [2019], Zhuang et al. [2020], Guo et al. [2021], Lei et al. [2021], Cao et al. [2022], Li et al. [2022] has been proposed. CDSR extends the SR framework by leveraging information across multiple domains, enabling a more holistic understanding of user preferences. By modeling interactions across diverse domains, CDSR systems can mitigate data sparsity and domain bias, leading to more accurate and comprehensive recommendations that reflect the user’s diverse interests and behaviors across different contexts.
Figure 1: (a) In the traditional CDSR framework, the input consists solely of item ID features. (b) In our LLM-EMF framework, we incorporate additional image and title information to complement the existing item features. Therefore, our method enriches item representations.
Early efforts in CDSR, such as $\pi$-Net Ma et al. [2019], focused on modeling item interaction sequences within a single domain and proposed transferring learned representations to other domains using gated transfer modules. These early models aimed to enhance cross-domain learning by reusing knowledge acquired from one domain to improve recommendation performance in other domains. However, these methods were primarily concerned with capturing intra-sequence relationships within individual domains, failing to address the complexities of inter-sequence relationships that span across multiple domains. As a result, while these approaches helped to some extent in alleviating the data sparsity and domain bias issues, they still faced challenges in fully integrating information from multiple domains in a meaningful way. Subsequent work, such as MIFN Ma et al. [2022], made strides by incorporating external knowledge graphs to strengthen the connections between domains. By leveraging knowledge graphs, these models could better align domain-specific information and enrich the representations of items and users across domains. Despite these improvements, the focus remained on enhancing the relationships within each individual domain, often neglecting the rich interactions between sequences across domains. This limitation further perpetuates the problem of domain bias, which continues to affect the overall performance of CDSR models. Consequently, there remains a significant need for techniques that can more effectively capture both intra- and inter-sequence relationships across domains to achieve optimal performance in real-world recommendation systems.
Moreover, existing CDSR methods fail to fully exploit the rich visual information associated with items. Visual cues play a critical role in shaping user preferences, as users often form initial impressions of items based on their visual appearance, even when textual descriptions are similar. To address this limitation, we propose LLM-Enhanced Multimodal Fusion for Cross-Domain Sequential Recommendation (LLM-EMF), a novel framework that leverages LLM knowledge to enhance textual information while integrating visual and textual embeddings to enrich item representations and model cross-domain user preferences. By leveraging the frozen CLIP model Radford et al. [2021], we generate aligned image and text embeddings, capturing both intra- and inter-sequence relationships. Our hierarchical attention mechanism jointly learns single-domain and cross-domain preferences, enabling effective modeling of user behavior across diverse domains.
Our key contributions are as follows:
  • 1.
    We incorporate LLM knowledge to enrich item representations by generating additional item-related text information, thus enhancing item representation.
  • 2.
    We propose the CDSR framework, which seamlessly integrates visual, textual, and item embeddings, thereby enriching item representations and boosting recommendation performance through multimodal fusion.
  • 3.
    We conduct comprehensive experiments across four reorganized e-commerce datasets, showing that our approach, LLM-EMF, outperforms state-of-the-art baselines on all evaluation metrics.
2 Related Work
2.1 Sequential Recommendation
Sequential recommendation Wang et al. [2019] serves as the fundamental framework for our approach, where user behavior is modeled as a sequence of time-sensitive items. The goal is to predict the likelihood of the next item in the sequence based on the user’s historical interactions. Initially, Markov chains were employed to capture the sequential dependencies within item sequences, as demonstrated in FPMC Rendle et al. [2010], where the assumption of a first-order Markov process enabled the modeling of user preferences over time. As the complexity of user behavior grew, deep learning techniques were introduced to capture more intricate patterns of sequential interactions. Recurrent neural networks (RNNs) Chung et al. [2014], Hochreiter & Schmidhuber [1997], convolutional neural networks (CNNs) LeCun et al. [1995], and attention mechanisms Vaswani et al. [2017] have since been integrated into recommender systems Hidasi et al. [2015], Zhou et al. [2019], Tang & Wang [2018], Kang & McAuley [2018a], Zhou et al. [2018], providing a more flexible and robust approach to modeling long-term dependencies and dynamic user preferences. These models can process sequential data more efficiently, making them highly effective in capturing the nuanced temporal evolution of user interests. In addition, methods like SURGE Chang et al. [2021] have leveraged metric learning techniques to reduce the complexity of item sequences, improving computational efficiency and enhancing the scalability of recommendation systems. This approach facilitates the efficient comparison of items in a lower-dimensional space, allowing for faster retrieval and better prediction accuracy in large-scale datasets.
More recently, works such as DFAR Lin et al. [2023] and DCN Lin et al. [2022] have sought to capture even deeper and more intricate relationships within sequential recommendation data. These models explore more sophisticated architectures, incorporating dual attention mechanisms and complex interaction learning to enhance the predictive power of the models, particularly in complex, noisy environments.
In this work, we extend the aforementioned techniques by applying cross-domain learning to sequential recommendation models. This novel approach facilitates the transfer of knowledge between multiple domains and LLM, allowing the model to leverage complementary information from different domains to improve recommendations. By doing so, we aim to enhance the generalization capabilities of sequential recommendation systems and make them more adaptable to a wide range of real-world applications.
2.2 Cross-Domain Recommendation
Cross-domain recommender systems Cantador et al. [2015] have emerged as a promising solution to address the challenges of sparse data and cold-start issues commonly encountered in sequential recommendation tasks. These systems leverage data from multiple domains to enrich user modeling, providing a more comprehensive understanding of user preferences. Early models in this area were largely grounded in single-domain recommendation frameworks, under the assumption that auxiliary user behaviors from different domains could be used to improve the modeling process in the target domain Singh & Gordon [2008]. This assumption led to the development of various methods that utilized data from multiple sources to better predict user preferences within a specific domain. A key technique in cross-domain recommendation is transfer learning Pan & Yang [2009], which involves transferring learned representations (such as user or item embeddings) from the source domain to the target domain. This transfer of knowledge helps enhance the recommendation process in the target domain by incorporating relevant information from other domains. Notable examples of transfer learning-based approaches include MiNet Ouyang et al. [2020], CoNet Hu et al. [2018], and itemCST Pan et al. [2010], among others. These methods have shown success in leveraging cross-domain knowledge to improve user-item interactions in the target domain. However, in practice, industrial platforms often aim to improve recommendation quality across multiple domains simultaneously, rather than focusing solely on enhancing the target domain while neglecting the source domain. This realization has led to the development of dual learning techniques Long et al. [2012], He et al. [2016], Hou et al. [2024], Li et al. [2025, 2024b, 2024a, 2024c], which enable mutual improvement between both the source and target domains. These methods aim to simultaneously enhance the performance in both domains by modeling their interactions, and have been successfully applied to cross-domain recommender systems Zhu et al. [2019], Li & Tuzhilin [2020]. By allowing both domains to benefit from each other’s knowledge, dual learning has opened up new possibilities for more robust and scalable recommendation systems. Despite these advancements, many existing methods fail to account for the role of visual perceptual features, which can provide valuable context for user preferences. Furthermore, many approaches do not integrate multimodal fusion with attention mechanisms, which can significantly enhance the ability of models to focus on relevant information from different sources. The integration of multimodal signals, such as images, text, and other media, with attention mechanisms can improve the model’s understanding of complex user preferences across domains. These innovations hold great potential for advancing the effectiveness of cross-domain recommendation systems.
2.3 LLM-based Recommendation System
Large Language Models (LLMs) have gained significant traction in recommender systems due to their ability to leverage textual features of items and improve recommendation performance Geng et al. [2022], Liu et al. [2023], Wu et al. [2023], Qiu et al. [2024, 2025b, 2025a]. Most existing LLM-based recommenders operate without fine-tuning, relying on pretrained knowledge to predict the next item Sun et al. [2023], Wang et al. [2023]. For example, CHAT-REC Gao et al. [2023] uses ChatGPT to understand user preferences and deliver interactive, explainable recommendations. Similarly, GPT4Rec Li et al. [2023] employs GPT-2 to generate hypothetical ”search queries” from user history, which are then processed by the BM25 search engine to retrieve relevant items. Another line of research focuses on fine-tuning LLMs for specific recommendation tasks. TALLRec Bao et al. [2023b], for instance, applies instruction-tuning to decide whether an item should be recommended, while BIGRec Bao et al. [2023a] combines LLM-generated recommendations with collaborative filtering through an ensemble approach. Additionally, RecInterpreter Yang et al. [2023] and LLaRA Liao et al. [2023] introduce the idea of treating ”sequential user behaviors” as a new modality for LLMs, aligning it with language representations. CoLLM Zhang et al. [2025] captures collaborative signals using an external model and integrates them into the LLM’s token embedding space, enabling the use of collaborative embeddings.
Despite the advancements in Sequential Recommendation (SR) through traditional or LLM-based methods, these approaches often adopt a narrow view of information formats, leading to underutilization of critical features such as the cold-start problem in Cross-Domain Sequential Recommendation (CDSR). Furthermore, while some studies like TALLRec Bao et al. [2023b], BIGRec Bao et al. [2023a], and LLM-Rec Tang et al. [2023] acknowledge the cross-domain potential of LLMs, there remains a gap in the development of LLM-based frameworks specifically tailored for CDSR. This paper aims to address these limitations by proposing a novel approach optimized for CDSR challenges.
Figure 2: The overview of our proposed LLM-EMF is as follows. Firstly, the Feature Preparation Process embeds all items from domains $X$ and $Y$ using a learnable item matrix for ID-based embeddings $E_{id}$, a frozen CLIP image encoder for image-based embeddings $E_{img}$, and a text encoder for text-based embeddings $E_{tex}$. Then, the input sequence $\mathcal{S}$ is forwarded to the Image Fusion Prediction Process. $\mathcal{S}$ consists of sub-sequences $S^{X}$, $S^{Y}$, and $S^{X+Y}$, which are processed through an embedding mapping layer to generate item ID-based, image-based, and text-based sequence embeddings. These embeddings are passed through multiple attention layers to capture both intra- and inter-sequence relationships, producing attention-aggregated embeddings. Finally, cosine similarity is applied between these aggregated embeddings and $E_{id}$, $E_{img}$, and $E_{tex}$ to predict the next item.
3 Methodology
3.1 Problem Formulation
[…]
Additionally, ${S^{X+Y}={\left[x_{1},y_{1},x_{2},\dots,x_{\left|S^{X}\right|},\dots,y_{\left|S^{Y}\right|}\right]}}$ represents the merged sequence by combining $S^{X}$ and $S^{Y}$. In general, the goal of the CDSR task is to predict the probabilities of candidate items across both domains and select the item with the highest probability as the next recommendation.
3.2 Overall framework
As illustrated in  fig. 2, the proposed LLM-EMF framework is designed to model cross-domain sequential preferences through a hierarchical architecture. The framework begins with item representation encoding, where items from domains $X$ and $Y$ are processed through parallel pathways to construct distinct representations. A learnable semantic embedding matrix $E_{id}$ is initialized to capture identity-specific features, while a frozen CLIP encoder generates perceptual embeddings $E_{img}$ and conceptual embeddings $E_{tex}$ from item images and textual descriptions, respectively. This multimodal encoding stage simulates the human capacity for integrating visual, semantic, and linguistic information.
The framework then proceeds to the preference simulation stage, where behavioral sequences $\mathcal{S}$ are processed. Each sequence $\mathcal{S}$ consists of three sub-sequences: $S^{X}$, $S^{Y}$, and $S^{X+Y}$, representing within-domain and cross-domain interactions. The embedding layer retrieves and restructures the multimodal embeddings ($E_{id}$, $E_{img}$, and $E_{tex}$) according to the sequence order, generating dynamic states for each sub-sequence. These states are further processed through hierarchical attention layers to model both intra-sequence and inter-sequence relationships, simulating the interplay between local working memory and global contextual integration.
Finally, the framework performs decision generation by comparing the attention-aggregated sequence embeddings with the original multimodal embeddings ($E_{id}$, $E_{img}$, and $E_{tex}$) using cosine similarity. This process mimics human decision-making by combining evidence from multiple sensory modalities to predict the next item in the sequence.
Figure 3: The prompt-and-generate pipeline of our method. In the showcase, the movie item is enhanced using the LLM with prompts to generate additional contextual information. As shown in the pipeline, the process begins by generating prompts, which are then input into the LLM. The output from the LLM consists of enhanced information, including key terms and a summary of the item, along with potential user interests. This enhanced information is subsequently used for the next step in text feature embedding.
3.3 Prompt with LLM
In the CDSR setting, we denote user-item interaction sequences as $X_{item}$ and their corresponding item metadata (e.g., titles and domains) as $T_{info}$. As illustrated in  fig. 3, for each of the items involved in the dataset, a pre-defined template is utilized as language commands for Deepseek-r1 Guo et al. [2025] to generate additional contextual knowledge. This approach ensures the relevance, diversity, and richness of item information, which is critical for enhancing the representation of items across multiple domains.
$P_{n}=\text{LLM}(commands),$ (1)
where $P_{n}$ represents the text prompt for generating enriched item representations, which is subsequently used as the input for enhancing item embeddings. For each item, we design a unified template as the input command for Deepseek-r1 Guo et al. [2025] to leverage the item’s title and description, generating additional contextual knowledge that enriches the item’s representation. This process requires only item-level metadata, making it scalable and efficient.
For instance, the template is designed as follows: “Generate additional contextual information for an item which is type of [domain] with [title]. Provide detailed insights and keywords about the item’s features. The output should include specific attributes, and potential users that could enhance the understanding of the item in a cross-domain recommendation context.
This approach not only augments the item’s metadata but also bridges the gap between domains by capturing domain-agnostic features and user preferences. By integrating the generated knowledge into the recommendation framework, we enable a more comprehensive and accurate modeling of user behavior across diverse domains. The enriched embeddings, combined with our hierarchical attention mechanism, facilitate the effective learning of both intra-domain and cross-domain relationships, ultimately improving the performance of the CDSR system.
3.4 Visual and Textual Feature Integration
We enhance user preferences by integrating multimodal features (image and textual) into the CDSR framework. This subsection provides a detailed explanation of feature preparation and sequence representation combined with image and textual features within a single domain, and applies this approach to domain $X$, domain $Y$, and domain $X+Y$.
1) Item Presentations Preparation with Image Feature: First, we prepare the feature representations for all items. As shown on the left in  fig. 2, we construct a learnable item matrix based on item IDs for all items in domain $X$ and $Y$, denoted as $E_{id}\in\mathbb{R}^{(|\mathcal{X}|+|\mathcal{Y}|)\times q}$. Here, $|\mathcal{X}|+|\mathcal{Y}|$ represents the total number of items from domains $X$ and $Y$, and $q$ is the learnable item embedding dimension. Simultaneously, we employ a pre-trained and frozen CLIP model to generate image embeddings for each item. These embeddings are used to form image matrix $E_{img}\in\mathbb{R}^{(|\mathcal{X}|+|\mathcal{Y}|)\times e}$, where $e$ is the image embedding dimension. These two matrices serve as the base for generating embeddings for sequences.
2) Item Presentations Preparation with Textual Feature: To further enrich the item representations, we leverage textual information from item titles using the same frozen CLIP model. For each item, the CLIP text encoder generates textual embeddings based on the item’s title, forming the textual matrix $E_{tex}\in\mathbb{R}^{(|\mathcal{X}|+|\mathcal{Y}|)\times e}$. This textual representation captures semantic information that complements the visual and ID-based features, enabling a more comprehensive understanding of user preferences.
3) Enhanced Similarity Score with Multimodal Features: We first generate a sequence representation to represent the user interaction sequence, which is then used to calculate the similarity with items for predicting the next item in the prediction process. As shown on the right in  fig. 2, we propose an embedding layer to process the input sequences $\mathcal{S}$, which consists of three sub-sequences $S^{X}$, $S^{Y}$, and $S^{X+Y}$. These sequences include item IDs, images, and textual data. Accordingly, the embedding layer produces item ID-based, image-based, and textual-based embeddings for each input sequence.
Once we obtain $E_{id}$, $E_{img}$, and $E_{tex}$, the item ID-based sequence embeddings $F_{id}\in\mathbb{R}^{|\mathcal{S}|\times q}$ for $\mathcal{S}$ can be generated by placing the appropriate embedding from $E_{id}$ in the order of the items in the sequence $\mathcal{S}$. Here, $|\mathcal{S}|$ denotes the total item number in $\mathcal{S}$. Similarly, the image-based sequence embeddings $F_{img}\in\mathbb{R}^{|\mathcal{S}|\times e}$ and textual-based sequence embeddings $F_{tex}\in\mathbb{R}^{|\mathcal{S}|\times e}$ are produced from $E_{img}$ and $E_{tex}$, respectively, based on the same sequence.
For simplicity, we take the image-based sequence embeddings $F_{img}$ as an example to illustrate how LLM-EMF obtains the next item prediction. We employ an Attention Layer to get the enhanced sequence image-based embeddings $H_{img}\in\mathbb{R}^{|\mathcal{S}|\times e}$ as follows:
$H_{img}=Attention(F_{img}),$ (2)
Afterward, we extract the last embedding vector $h_{img}\in\mathbb{R}^{1\times e}$ from $H_{img}$ as the sequence representation, which serves as the attention-aggregated embedding of the sequence. This vector captures the user preferences, specifically focusing on the most recent interaction within the sequence. Then, $h_{img}$ is compared against $E_{img}$ using cosine similarity. In this way, the alignment between user preferences and the embedding of each item across all domains is assessed. The similarity score is computed as follows:
$Sim(h_{img},E_{img})=\frac{h_{img}\cdot E_{img}^{T}}{\|h_{img}\|\|E_{img}^{T}\|},$ (3)
where $T$ denotes the transpose function. Here, a higher similarity score indicates that the sequence preference is more aligned with this item. In the same manner, we can obtain the item ID-based sequence representative vector $h_{id}\in\mathbb{R}^{1\times q}$ and textual-based sequence representative vector $h_{tex}\in\mathbb{R}^{1\times e}$, along with their respective similarity scores. In the next section, we will introduce how the multiple attention layer is used to fuse ID-based, image-based, and textual-based predictions for CDSR.
Figure 4: The process of transforming a user sequence into a representation sequence. Initially, the user’s sequence of interactions is converted into an embedded representation. This sequence is then passed through an attention layer to capture both intra- and inter-sequence relationships. Finally, the attention-aggregated sequence representations are compared with item embeddings, and the item with the highest similarity score is selected as the predicted next item.
3.5 Multiple Attention Mechanisms
In traditional sequential recommendation methods Hidasi et al. [2015], Kang & McAuley [2018b], behavioral sequences from domains $X$ and $Y$ are often processed indiscriminately, leading to potential dominance by the domain with higher behavioral frequency. This imbalance can skew the representation of user preferences. To address this issue, we propose a hierarchical attentional framework that separately processes sequences $S^{X}$, $S^{Y}$, and $S^{X+Y}$, ensuring balanced representation across domains. Our approach integrates item identity (ID), visual features, and textual embeddings to model user preferences more comprehensively. As illustrated in  fig. 4, to model the interactions within each sequence more effectively, we incorporate a self-attention mechanism with query, key, and value (QKV) pairs. The attention mechanism allows the model to focus on relevant parts of the sequence when making predictions, capturing both local and global dependencies within the sequence.
The attention mechanism computes the output as a weighted sum of the values, where the weights are determined by the similarity between the query and the keys:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V,$ (4)
where $Q$, $K$, and $V$ are the query, key, and value matrices, respectively, and $d_{k}$ is the dimension of the key vectors. This mechanism is applied within each domain’s sequence processing to capture both intra-domain and cross-domain relationships more effectively.
The framework generates nine sequence representations through multiple attention layers: $h^{X}_{id}$, $h^{X}_{img}$, $h^{X}_{tex}$, $h^{Y}_{id}$, $h^{Y}_{img}$, $h^{Y}_{tex}$, $h^{X+Y}_{id}$, $h^{X+Y}_{img}$, and $h^{X+Y}_{tex}$. These representations capture both intra-domain and cross-domain user preferences. For domain $X$, the prediction probability for the next item $x_{t+1}$ in sequence $S^{X}$ is computed as follows:
$\small\mathrm{P}^{X}_{id}(x_{t+1}\mid\mathcal{S})=\text{softmax}\left(\text{Sim}(h_{id}^{X},E_{id}^{X})\right),\quad x_{t}\in S^{X},$ (5)
where $x_{t+1}$ denotes the predicted next item, $h_{id}^{X}$ is the ID-based sequence representation, and $E_{id}^{X}$ is the ID embedding matrix for domain $X$. Similarly, the prediction probabilities based on visual and textual features are calculated as:
$\small\mathrm{P}^{X}_{img}(x_{t+1}\mid\mathcal{S})=\text{softmax}\left(\text{Sim}(h_{img}^{X},E_{img}^{X})\right),\quad x_{t}\in S^{X},$ (6)
$\small\mathrm{P}^{X}_{tex}(x_{t+1}\mid\mathcal{S})=\text{softmax}\left(\text{Sim}(h_{tex}^{X},E_{tex}^{X})\right),\quad x_{t}\in S^{X},$ (7)
where $h_{img}^{X}$, $E_{img}^{X}$, $h_{tex}^{X}$, and $E_{tex}^{X}$ represent the image-based sequence representation, image embedding matrix, text-based sequence representation, and text embedding matrix, respectively. The final prediction probability for the next item combines ID, image, and text-based predictions:
$\mathrm{P}^{X}(x_{t+1}\mid\mathcal{S})$ $=\alpha\mathrm{P}^{X}_{id}(x_{t+1}\mid\mathcal{S})$
$\quad+\beta\mathrm{P}^{X}_{img}(x_{t+1}\mid\mathcal{S})$
$\quad+(1-\alpha-\beta)\mathrm{P}^{X}_{tex}(x_{t+1}\mid\mathcal{S}),$ (8)
where $\alpha$ and $\beta$ are weighting factors that balance the contributions of ID, visual, and textual features. This formulation captures structural, perceptual, and semantic aspects of user preferences.
To optimize the framework, we minimize the following loss function for domain $X$:
$\small\mathcal{L}^{X}=\sum_{x_{t}\in S^{X}}-\log\mathrm{P}^{X}(x_{t+1}\mid\mathcal{S}).$ (9)
The same method is applied to domains $Y$ and $X+Y$, yielding loss terms $\mathcal{L}^{Y}$ and $\mathcal{L}^{X+Y}$. The final loss is a weighted combination of these terms:
$\mathcal{L}=\mathcal{L}^{X}+\lambda_{1}\mathcal{L}^{Y}+\lambda_{2}\mathcal{L}^{X+Y},$ (10)
where $\lambda_{1}$ and $\lambda_{2}$ control the relative importance of each domain.
During evaluation, we aggregate predictions from all domains to recommend the next item. Let $X$ be the target domain and $Y$ the source domain. The combined prediction score for item $x_{i}$ is computed as:
$\small\mathrm{P}(x_{i}|S)=\mathcal{P}^{X}(x_{i}|S)+\lambda_{1}\mathcal{P}^{Y}(x_{i}|S)+\lambda_{2}\mathcal{P}^{X+Y}(x_{i}|S).$ (11)
The recommended item is selected by maximizing the prediction score within the target domain $X$:
$\mathrm{argmax}_{x_{i}\in\mathcal{X}}\mathrm{P}(x_{i}|S).$ (12)
This process allows the model to incorporate cross-domain information effectively, leveraging the predictions from both the target domain and the source domain. By weighting the contributions from each domain, the model can refine its recommendation, taking into account the richer information that might come from the source domain while focusing on the target domain for the final prediction.
The weighting factors $\lambda_{1}$ and $\lambda_{2}$ allow the model to control how much influence the source domain ($Y$) and the combined domain ($X+Y$) have on the recommendation for the target domain. This flexibility in adjusting the weights is crucial for handling different types of datasets or domain-specific characteristics, ensuring that the model adapts to varying degrees of relevance between domains.
Finally, by selecting the item with the highest prediction score from the target domain, the model aims to provide the most contextually appropriate recommendation, which is critical for real-world recommendation systems that need to balance both accuracy and relevance across multiple domains.
Table 1: Statistics of two CDSR scenarios.
Scenarios #Items #Train #Valid #Test Avg.length
Food 29,207 34,117 2,722 2,747 9.91
Kitchen 34,886 5,451 5,659
Movie 36,845 58,515 2,032 1,978 11.98
Book 63,937 5,612 5,730
Table 2: Experimental results (%) on the Food-Kitchen scenario.
Model (Food-Kitchen) MRR NDCG MRR NDCG
(Food) @5 @10 (Kitchen) @5 @10
ItemKNN Sarwar et al. [2001] 3.92 3.51 3.97 1.89 1.28 1.75
NCF-MLP He et al. [2017b] 4.49 3.94 4.51 2.18 1.57 2.03
BPRMF Rendle et al. [2012] 4.10 3.55 4.03 2.01 1.45 1.85
$\pi$-Net Ma et al. [2019] 7.68 7.32 8.13 3.53 2.98 3.73
GRU4Rec Hidasi et al. [2015] 5.79 5.48 6.13 3.06 2.55 3.10
SASRec Kang & McAuley [2018b] 7.30 6.90 7.79 3.79 3.35 3.93
SR-GNN Wu et al. [2019] 7.84 7.58 8.35 4.01 3.47 4.13
PSJNet Sun et al. [2022] 8.33 8.07 8.77 4.10 3.68 4.32
MIFN Ma et al. [2022] 8.55 8.28 9.01 4.09 3.57 4.29
Tri-CDR Ma et al. [2024] 8.35 8.18 8.89 4.29 3.63 4.33
Ours LLM-EMF 9.24 9.05 10.21 5.13 4.84 5.43
Table 3: Experimental results (%) on the Movie-Book scenario.
Model (Movie-Book) MRR NDCG MRR NDCG
(Movie) @5 @10 (Book) @5 @10
ItemKNN Sarwar et al. [2001] 2.92 2.17 2.88 1.26 0.86 1.10
NCF-MLP He et al. [2017b] 3.05 2.26 2.96 1.43 1.06 1.26
BPRMF Rendle et al. [2012] 2.96 2.18 2.80 1.27 0.85 1.17
$\pi$-Net Ma et al. [2019] 4.16 3.72 4.17 2.17 1.84 2.03
GRU4Rec Hidasi et al. [2015] 3.83 3.14 3.73 1.68 1.34 1.52
SASRec Kang & McAuley [2018b] 3.79 3.23 3.69 1.81 1.41 1.71
SR-GNN Wu et al. [2019] 3.85 3.27 3.78 1.78 1.40 1.66
PSJNet Sun et al. [2022] 4.63 4.06 4.76 2.44 2.07 2.35
MIFN Ma et al. [2022] 5.05 4.21 5.20 2.51 2.12 2.31
Tri-CDR Ma et al. [2024] 5.15 4.62 5.05 2.32 2.08 2.22
Ours LLM-EMF 6.32 5.35 5.78 2.86 2.53 2.79
4 Experiments
In this section, we introduce the experimental setup, providing detailed descriptions of the dataset, evaluation metrics, and implementation specifics. We then compare our method with state-of-the-art approaches on the Amazon dataset Wei et al. [2021]. Finally, we conduct ablation studies to validate the effectiveness of the proposed method.
Dataset and Evaluated Metric. Our experiments are conducted on the Amazon dataset Wei et al. [2021], which serves as the foundation for constructing cross-domain sequential recommendation (CDSR) scenarios. The dataset offers a rich collection of user-item interactions across a variety of domains, making it ideal for modeling cross-domain behaviors. Building on previous works Ma et al. [2019, 2022], we select four distinct domains to create two CDSR scenarios for our experiments: “Food-Kitchen” and “Movie-Book.” These domains were chosen to reflect real-world applications where users frequently engage with multiple categories, allowing us to investigate the efficacy of cross-domain knowledge transfer in sequential recommendation settings. The process of preparing the data begins by extracting users who have interacted with items in both domains. To ensure meaningful analysis, we filter out users and items with fewer than 10 interactions, focusing only on those with sufficient behavioral data. This filtering step is crucial for ensuring the robustness of our results, as sparse data can lead to unreliable recommendations. Furthermore, in order to meet the sequential constraints necessary for our study, we retain cross-domain interaction sequences that contain at least three items from each domain. This ensures that each sequence provides enough context for learning temporal dependencies and domain transitions. For the training/validation/test split, we adopt a temporal strategy to preserve the sequential nature of the data. Specifically, we divide the latest interaction sequences into validation and test sets, while the remaining sequences are allocated to the training set. This approach helps simulate a real-world scenario, where past user behavior informs future recommendations. The statistics of our refined datasets for the CDSR scenarios, including the number of users, items, and interactions in each domain, are summarized in  table 1.
To evaluate the performance of our model, we adopt two widely used metrics in recommendation systems: Mean Reciprocal Rank (MRR) Voorhees [1999] and Normalized Discounted Cumulative Gain (NDCG@5, 10) Järvelin & Kekäläinen [2002]. MRR is a metric that assesses the ranking quality by calculating the average of the reciprocal ranks of the first relevant item in each user’s recommendation list. NDCG, on the other hand, takes into account both the position and relevance of the recommended items, making it particularly suitable for evaluating ranked lists. These metrics are standard in the evaluation of recommendation systems, providing a comprehensive measure of recommendation quality. The results obtained from these metrics allow us to compare our model’s performance across different CDSR scenarios and demonstrate its effectiveness in capturing cross-domain interactions.
Figure 5: Impact of Hyperparameters on Final Performance on Movie Data
Implementation Details. For fair comparisons, we adopt the same hyperparameter settings as in previous works Ma et al. [2022]. We conduct experiments with various hyperparameter settings, as shown in Figure 5. Specifically, we set the learnable embedding size $q$ to 256, the CLIP-based image and text embedding size $e$ to 512, and the mini-batch size to 256, with a dropout rate of 0.3. The values of $\lambda_{1}$ and $\lambda_{2}$ are set to 0.3 and 0.1, respectively. The $L_{2}$ regularization coefficient is selected from $\{0.00001,0.00005,0.0001,0.0005,0.001\}$, and the learning rate is chosen from $\{0.0001,0.0005,0.001,0.005,0.01\}$. Training is conducted for 100 epochs using an NVIDIA 4090 GPU, with the Adam optimizer Kingma & Ba [2015] to update all parameters. To ensure robustness and reproducibility, we perform a grid search over the hyperparameter space, evaluating the model’s performance on a validation set at each step. The selected hyperparameters are those that yield the best validation performance. Additionally, we employ early stopping with a patience of 10 epochs to prevent overfitting, monitoring the validation loss during training. This approach ensures that the model generalizes well to unseen data while maintaining computational efficiency. The use of the Adam optimizer Kingma & Ba [2015] is particularly advantageous due to its adaptive learning rate mechanism, which helps stabilize training and accelerate convergence. Furthermore, the NVIDIA 4090 GPU provides the necessary computational power to handle the large-scale datasets and complex model architectures efficiently. These settings collectively ensure that our experiments are conducted under optimal conditions, enabling meaningful comparisons with state-of-the-art methods in the field.
4.1 Performance Comparisons
We evaluate our proposed LLM-EMF method against a comprehensive set of state-of-the-art (SOTA) approaches on the “Food-Kitchen” and “Movie-Book” CDSR scenarios. The experimental results are summarized in  table 2 and  table 3, respectively. Our approach consistently outperforms all existing methods across both scenarios, demonstrating its effectiveness in handling cross-domain recommendation tasks. In the Food-Kitchen scenario, our LLM-EMF model achieves the highest scores across all evaluation metrics. Specifically, it attains an MRR of 9.24% for Food and 5.13% for Kitchen, significantly outperforming the closest competitor, MIFN, which achieves MRR scores of 8.55% and 4.09%, respectively. In terms of NDCG@5 and NDCG@10, LLM-EMF also leads with 9.05% and 10.21% for Food, and 4.84% and 5.43% for Kitchen. The performance of traditional methods such as ItemKNN, NCF-MLP, and BPRMF is notably lower, highlighting the limitations of these approaches in capturing complex user-item interactions in cross-domain settings. More advanced models like $\pi$-Net, GRU4Rec, and SASRec show improved performance but still fall short of LLM-EMF. This suggests that our model’s ability to leverage large language models (LLMs) for enhanced feature extraction and cross-domain knowledge transfer is a key factor in its superior performance. In the Movie-Book scenario, LLM-EMF again demonstrates its superiority, achieving MRR scores of 6.32% for Movie and 2.86% for Book. These results are significantly higher than those of the next best model, MIFN, which achieves MRR scores of 5.05% and 2.51%, respectively. Similarly, in terms of NDCG@5 and NDCG@10, LLM-EMF leads with 5.35% and 5.78% for Movie, and 2.53% and 2.79% for Book.
Table 4: Ablation study on the movie dataset
original-framework Textual Fusion LLM-enhanced Visual Fusion MRR 5.03 5.88 6.01 6.08 6.15 6.32
4.2 Ablation Studies
We conducted an ablation study to systematically evaluate the contributions of our proposed components: Textual Fusion, LLM-enhanced, and Visual Fusion. As illustrated in  table 4, we first established a baseline using the original framework without any additional modules, achieving an MRR of 5.03. The introduction of Textual Fusion alone resulted in a performance gain of 0.85 points, increasing the MRR to 5.88, which highlights the importance of leveraging textual information for cross-domain recommendation tasks. When the LLM-enhanced module was added alongside Textual Fusion, the MRR further improved to 6.01, demonstrating the effectiveness of leveraging large language models to enrich textual representations. Notably, the inclusion of Visual Fusion alone significantly boosted the MRR to 6.08, underscoring the critical role of visual signals in enhancing the model’s understanding of user preferences. The combination of Textual Fusion and Visual Fusion achieved an MRR of 6.15, indicating that textual and visual information complement each other in capturing cross-domain user behavior. Finally, the full framework, incorporating Textual Fusion, LLM-enhanced, and Visual Fusion, achieved the highest MRR of 6.32. This result demonstrates the cumulative impact of integrating multiple modalities and advanced language models for improving recommendation performance.
5 Conclusion
In this study, we propose a novel framework, named LLM-EMF, for cross-domain sequential recommendation. Unlike traditional methods that focus solely on temporal patterns within user interaction sequences, our approach leverages LLM knowledge to enrich textual information while also utilizing the powerful visual and textual understanding capabilities of the CLIP model to integrate multimodal representations. Specifically, we design prompts to combine the title and domain information, which are then used to generate augmented text. Both the images and augmented text are processed through CLIP to generate embeddings, which are subsequently fused with a learnable item matrix to enhance item representations. By combining these innovations, our LLM-EMF framework achieves state-of-the-art performance in CDSR, demonstrating the effectiveness of integrating multimodal signals and LLM knowledge in sequential recommendation tasks.
6 Acknowledgements
This work was supported by the National Natural Science Foundation of China (No. 62471405, 62331003, 62301451), Suzhou Basic Research Program (SYG202316) and XJTLU REF-22-01-010, XJTLU AI University Research Centre, Jiangsu Province Engineering Research Centre of Data Science and Cognitive Computation at XJTLU and SIP AI innovation platform (YZCXPT2022103).