### v2

# LLM-Enhanced Multimodal Fusion for Cross-Domain Sequential Recommendation

Abstract
Cross-Domain Sequential Recommendation (CDSR) predicts user behavior by leveraging historical interactions across multiple domains, capturing both intra- and inter-sequence item relationships. To further enhance the value of visual and textual data, we propose LLM-EMF, an innovative approach that incorporates Large Language Models(LLM) to enrich textual data and boosts recommendation performance by merging visual and textual information. Additionally, a multi-attention mechanism is designed to jointly learn single-domain and cross-domain preferences, effectively capturing complex user interests. Evaluations on four e-commerce datasets demonstrate that LLM-EMF outperforms existing methods in modeling cross-domain user preferences, highlighting the advantages of multimodal integration in sequential recommendation systems.
keywords:
Large Language Models, CDSR, CLIP-based Embeddings
1 Introduction
Sequential Recommendation (SR) has become a widely adopted approach for modeling dynamic user preferences by analyzing historical interaction sequences [1, 2, 3, 4]. While SR effectively predicts the next item within a single domain, it suffers from data sparsity and domain bias, where limited interaction data and overfitting to domain-specific patterns hinder generalization. Cross-Domain Sequential Recommendation (CDSR) [5, 6, 7, 8, 9, 10, 11] addresses these challenges by leveraging user behavior across multiple domains, enabling knowledge transfer and a more holistic modeling of user interests. However, existing CDSR methods still face three key limitations: (i) they primarily focus on intra-domain dependencies, underexploring inter-domain relationships; (ii) they insufficiently utilize multimodal information such as images and text; and (iii) even LLM-based approaches [12] rarely consider domain imbalance in cross-domain preference modeling. The difference between traditional CDSR and our approach is shown in Fig. 1.
Figure 1: (a) Traditional CDSR uses only item ID features. (b) Our LLM-EMF incorporates image and title information, enriching item representations.
To address these gaps, we propose LLM-Enhanced Multimodal Fusion for Cross-Domain Sequential Recommendation (LLM-EMF), a novel framework that integrates prompt-designed Large Language Model (LLM) augmentation, multimodal fusion, and domain-balanced hierarchical attention. Unlike LLMRec [12], which augments textual features without explicitly mitigating domain imbalance, LLM-EMF generates domain-agnostic semantic attributes via LLM prompts to enhance cross-domain alignment. Compared with IFCDSR [13], which fuses image and ID embeddings but lacks explicit textual enrichment, our framework unifies visual, textual, and ID-based embeddings. In contrast to MAN [14], which applies mixed attention without balancing domain contributions, our hierarchical attention mechanism explicitly regulates domain influence, preventing high-frequency domains from dominating recommendations. To the best of our knowledge, LLM-EMF is the first CDSR framework that systematically integrates LLM-generated textual knowledge with multimodal fusion and explicit domain balance.
Our key contributions are as follows:
  • 1.
    We design a prompt-based LLM augmentation strategy to generate additional domain-agnostic textual attributes, improving semantic alignment between domains.
  • 2.
    We propose a multimodal fusion framework that jointly models visual, textual, and ID embeddings via a hierarchical attention mechanism, capturing both intra- and inter-domain preferences while mitigating domain imbalance.
  • 3.
    We conduct extensive experiments on four reorganized e-commerce datasets, demonstrating that LLM-EMF consistently outperforms recent state-of-the-art baselines [13, 14] across all evaluation metrics.
2 Related Work
2.1 Sequential Recommendation
Sequential recommendation [15, 16, 17, 18] models user behavior as a sequence of time-sensitive items, aiming to predict the next item based on historical interactions. Early approaches, such as FPMC [19], used first-order Markov chains to model temporal dependencies. With growing behavioral complexity, deep learning models—RNNs [20, 21], CNNs [22], and attention mechanisms [23, 24, 25, 26, 27, 28]—were integrated into recommender systems [29, 30, 31, 32, 33], capturing long-term dependencies and dynamic preferences. Methods like SURGE [34] further reduced sequence complexity via metric learning for efficient large-scale retrieval. More recently, works such as DFAR [35] and DCN [36] have sought to capture deeper and more intricate relationships within sequential recommendation data using dual attention mechanisms and complex interaction modeling.
Our work extends these techniques with cross-domain learning, enabling knowledge transfer between domains and LLMs to leverage complementary signals, thereby improving generalization and adaptability.
2.2 Cross-Domain Sequential Recommendation
Unlike single-domain sequential recommendation, which models behavior within one domain, cross-domain recommender systems [37, 38] address sparse data and cold-start issues by leveraging user behavior from multiple domains. Early methods [39] assumed auxiliary behaviors from other domains could enhance the target domain. Transfer learning [40] approaches (e.g., MiNet [41], CoNet [42], itemCST [43]) transfer learned representations to boost target-domain recommendations.
Industrial settings often require improving performance in all domains simultaneously, motivating dual learning [44, 45, 46, 47, 48, 49, 50, 51, 52], as in DTCDR and DDTCDR, to enable mutual enhancement via domain interaction modeling. However, many approaches overlook visual perceptual features and lack multimodal fusion with attention mechanisms, which can better focus on relevant signals from images, text, and other media, enriching user preference modeling across domains.
2.3 LLM-based Recommendation System
Large Language Models (LLMs) have gained significant traction in recommender systems due to their ability to leverage textual features of items and improve recommendation performance [53, 54, 55, 56, 57, 58]. Most existing LLM-based recommenders operate without fine-tuning, relying on pretrained knowledge to predict the next item [59, 60]. For example, CHAT-REC [61] uses ChatGPT to understand user preferences and deliver interactive, explainable recommendations. Similarly, GPT4Rec [62] employs GPT-2 to generate hypothetical ”search queries” from user history, which are then processed by the BM25 search engine to retrieve relevant items. Another line of research focuses on fine-tuning LLMs for specific recommendation tasks. TALLRec [63], for instance, applies instruction-tuning to decide whether an item should be recommended, while BIGRec [64] combines LLM-generated recommendations with collaborative filtering through an ensemble approach. Additionally, RecInterpreter [65] and LLaRA [66] treat ”sequential user behaviors” as a new modality for LLMs, aligning it with language representations. CoLLM [67] captures collaborative signals using an external model and integrates them into the LLM’s token embedding space, enabling the use of collaborative embeddings.
While some studies like TALLRec [63], BIGRec [64], and LLM-Rec [68] acknowledge the cross-domain potential of LLMs, there remains a gap in frameworks explicitly designed for CDSR that combine LLM-based textual enrichment with multimodal fusion and domain-balanced attention.
Compared with recent representative methods, our work addresses several gaps. First, unlike LLMRec [12], which augments item text but does not explicitly mitigate domain imbalance, our framework applies a prompt-designed LLM augmentation strategy to generate domain-agnostic attributes that improve cross-domain semantic alignment. Second, in contrast to IFCDSR [13], which fuses image and ID embeddings without enriched textual features, we unify visual, textual, and ID representations within a single multimodal fusion framework. Finally, whereas MAN [14] employs mixed attention without controlling domain contributions, our hierarchical attention mechanism explicitly balances the influence of each domain, preventing high-frequency domains from dominating recommendations.
Figure 2: Overview of the proposed LLM-EMF framework. The Feature Preparation module generates ID-, image-, and text-based embeddings from domains $X$ and $Y$ using a learnable ID matrix, a frozen CLIP image encoder, and a text encoder. These embeddings are processed through multi-layer attention to model intra- and inter-sequence relationships, and cosine similarity with the embedding matrices is used for next-item prediction.
3 Methodology
3.1 Problem Formulation
In the CDSR task, user interaction sequences occur in domain $X$ and domain $Y$, with their item sets denoted as $\mathcal{X}$ and $\mathcal{Y}$, respectively. Let $\mathcal{S}$ represent the overall interaction sequence of the user in chronological order, which consists of three sub-sequences $(S^{X},S^{Y},S^{X+Y})\in\mathcal{S}$}. Specifically, let $S^{X}=\left[x_{1},x_{2},\dots,x_{\left|S^{X}\right|}\right],x\in\mathcal{X}$ and $S^{Y}=\left[y_{1},y_{2},\dots,y_{\left|S^{Y}\right|}\right],y\in\mathcal{Y}$ represent the interaction sequences within each domain, where $\left|\cdot\right|$ denotes the total number of items in each sequence.
Additionally, $S^{X+Y}={\left[x_{1},y_{1},x_{2},\dots,x_{\left|S^{X}\right|},\dots,y_{\left|S^{Y}\right|}\right]}$} represents the merged sequence by combining $S^{X}$ and $S^{Y}$. In general, the goal of the CDSR task is to predict the probabilities of candidate items across both domains and select the item with the highest probability as the next recommendation.
3.2 Overall framework
As illustrated in Fig. 2, the proposed LLM-EMF framework is designed to model cross-domain sequential preferences through a hierarchical architecture. The framework begins with item representation encoding, where items from domains $X$ and $Y$ are processed through parallel pathways to construct distinct representations. A learnable semantic embedding matrix $E_{id}$ is initialized to capture identity-specific features, while a frozen CLIP encoder generates perceptual embeddings $E_{img}$ and conceptual embeddings $E_{tex}$ from item images and textual descriptions, respectively. This multimodal encoding stage simulates the human capacity for integrating visual, semantic, and linguistic information.
The framework then proceeds to the preference simulation stage, where behavioral sequences $\mathcal{S}$ are processed. Each sequence $\mathcal{S}$ consists of three sub-sequences: $S^{X}$, $S^{Y}$, and $S^{X+Y}$, representing within-domain and cross-domain interactions. The embedding layer retrieves and restructures the multimodal embeddings ($E_{id}$, $E_{img}$, and $E_{tex}$) according to the sequence order, generating dynamic states for each sub-sequence. These states are further processed through hierarchical attention layers to model both intra-sequence and inter-sequence relationships, simulating the interplay between local working memory and global contextual integration.
Finally, the framework performs decision generation by comparing the sequence embeddings with the original multimodal embeddings ($E_{id}$, $E_{img}$, and $E_{tex}$) using cosine similarity. This process mimics human decision-making by combining evidence from multiple sensory modalities to predict the next item in the sequence.
Figure 3: The prompt-and-generate pipeline of our method. In the showcase, the movie item is enhanced using the LLM with prompts to generate additional contextual information. As shown in the pipeline, the process begins by generating prompts, which are then input into the LLM. The output from the LLM consists of enhanced information, including key terms and a summary of the item, along with potential user interests. This enhanced information is subsequently used for the next step in text feature embedding.
3.3 Prompt with LLM
In the CDSR setting, we denote user–item interaction sequences as $X_{item}$ and their corresponding item metadata (e.g., titles and domains) as $T_{info}$. As illustrated in Fig. 3, for each item in the dataset, a pre-defined template is used as a language command for Deepseek-r1 [69] to generate additional contextual knowledge. This approach ensures the relevance, diversity, and richness of item information, which is critical for enhancing item representations across multiple domains.
$P_{n}=\text{LLM}(commands),$ (1)
where $P_{n}$ represents the generated text prompt for enriching item representations, which is subsequently used as input for enhancing item embeddings. For each item, we design a unified template as the input command for Deepseek-r1 [69], leveraging the item’s title and description to generate additional contextual knowledge that further enriches the representation. This process requires only item-level metadata, making it both scalable and efficient.
For instance, the template is designed as follows: “Generate additional contextual information for an item of type [domain] with [title]. Provide detailed insights and keywords about the item’s features. The output should include specific attributes and potential users that could enhance the understanding of the item in a cross-domain recommendation context.”
This approach not only augments the item’s metadata but also bridges gaps between domains by capturing domain-agnostic features and user preferences. By integrating the generated knowledge into the recommendation framework, we enable more comprehensive and accurate modeling of user behavior across diverse domains. The enriched embeddings, combined with our hierarchical attention mechanism, facilitate effective learning of both intra-domain and cross-domain relationships, ultimately improving the performance of the CDSR system.
3.4 Visual and Textual Feature Integration
We enhance user preferences by integrating multimodal features (image and textual) into the CDSR framework. This subsection provides a detailed explanation of feature preparation and sequence representation combined with image and textual features within a single domain, and applies this approach to domain $X$, domain $Y$, and domain $X+Y$.
1) Item Presentations Preparation with Image Feature: First, we prepare the feature representations for all items. As shown on the left in Fig. 2, we construct a learnable item matrix based on item IDs for all items in domain $X$ and $Y$, denoted as $E_{id}\in\mathbb{R}^{(|\mathcal{X}|+|\mathcal{Y}|)\times q}$. Here, $|\mathcal{X}|+|\mathcal{Y}|$ represents the total number of items from domains $X$ and $Y$, and $q$ is the learnable item embedding dimension. Simultaneously, we employ a pre-trained and frozen CLIP model to generate image embeddings for each item. These embeddings are used to form image matrix $E_{img}\in\mathbb{R}^{(|\mathcal{X}|+|\mathcal{Y}|)\times e}$, where $e$ is the image embedding dimension. These two matrices serve as the base for generating embeddings for sequences.
2) Item Presentations Preparation with Textual Feature: To further enrich the item representations, we leverage textual information from item titles using the same frozen CLIP model. For each item, the CLIP text encoder generates textual embeddings based on the item’s title, forming the textual matrix $E_{tex}\in\mathbb{R}^{(|\mathcal{X}|+|\mathcal{Y}|)\times e}$. This textual representation captures semantic information that complements the visual and ID-based features, enabling a more comprehensive understanding of user preferences.
3) Enhanced Similarity Score with Multimodal Features: We first generate a sequence representation to represent the user interaction sequence, which is then used to calculate the similarity with items for predicting the next item in the prediction process. As shown on the right in Fig. 2, we propose an embedding layer to process the input sequences $\mathcal{S}$, which consists of three sub-sequences $S^{X}$, $S^{Y}$, and $S^{X+Y}$. These sequences include item IDs, images, and textual data. Accordingly, the embedding layer produces item ID-based, image-based, and textual-based embeddings for each input sequence.
Once we obtain $E_{id}$, $E_{img}$, and $E_{tex}$, the item ID-based sequence embeddings $F_{id}\in\mathbb{R}^{|\mathcal{S}|\times q}$ for $\mathcal{S}$ can be generated by placing the appropriate embedding from $E_{id}$ in the order of the items in the sequence $\mathcal{S}$. Here, $|\mathcal{S}|$ denotes the total item number in $\mathcal{S}$. Similarly, the image-based sequence embeddings $F_{img}\in\mathbb{R}^{|\mathcal{S}|\times e}$ and textual-based sequence embeddings $F_{tex}\in\mathbb{R}^{|\mathcal{S}|\times e}$ are produced from $E_{img}$ and $E_{tex}$, respectively, based on the same sequence.
For simplicity, we take the image-based sequence embeddings $F_{img}$ as an example to illustrate how LLM-EMF obtains the next item prediction. We employ an Attention Layer to obtain the enhanced sequence of image-based embeddings $H_{img}\in\mathbb{R}^{|\mathcal{S}|\times e}$, which captures the most relevant visual features from the input sequence for improved representation learning, as follows:
$H_{img}=Attention(F_{img}).$ (2)
Afterward, we extract the last embedding vector $h_{img}\in\mathbb{R}^{1\times e}$ from $H_{img}$ as the sequence representation, which serves as the attention-aggregated embedding of the sequence. This vector captures the user preferences, specifically focusing on the most recent interaction within the sequence. Then, $h_{img}$ is compared against $E_{img}$ using cosine similarity. In this way, the alignment between user preferences and the embedding of each item across all domains is assessed. The similarity score is computed as follows:
$Sim(h_{img},E_{img})=\frac{h_{img}\cdot E_{img}^{T}}{\|h_{img}\|\|E_{img}^{T}\|},$ (3)
where $T$ denotes the transpose function. Here, a higher similarity score indicates that the sequence preference is more aligned with this item. In the same manner, we can obtain the item ID-based sequence representative vector $h_{id}\in\mathbb{R}^{1\times q}$ and textual-based sequence representative vector $h_{tex}\in\mathbb{R}^{1\times e}$, along with their respective similarity scores. In the next section, we will introduce how the multiple attention layer is used to fuse ID-based, image-based, and textual-based predictions for CDSR.
Figure 4: The process of transforming a user sequence into a representation sequence. Initially, the user’s sequence of interactions is converted into an embedded representation. This sequence is then passed through an attention layer to capture both intra- and inter-sequence relationships. Finally, the attention-aggregated sequence representations are compared with item embeddings, and the item with the highest similarity score is selected as the predicted next item.
3.5 Multiple Attention Mechanisms
In traditional sequential recommendation methods [29, 70], behavioral sequences from domains $X$ and $Y$ are often processed indiscriminately, leading to potential dominance by the domain with higher behavioral frequency. This imbalance can skew the representation of user preferences. To address this issue, we propose a hierarchical attentional framework that separately processes sequences $S^{X}$, $S^{Y}$, and $S^{X+Y}$, ensuring balanced representation across domains. Our approach integrates item identity (ID), visual features, and textual embeddings to model user preferences more comprehensively. As illustrated in Fig. 4, to model the interactions within each sequence more effectively, we incorporate a self-attention mechanism with query, key, and value (QKV) pairs. The attention mechanism allows the model to focus on relevant parts of the sequence when making predictions, capturing both local and global dependencies within the sequence.
The attention mechanism computes the output as a weighted sum of the values, where the weights are determined by the similarity between the query and the keys:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V,$ (4)
where $Q$, $K$, and $V$ are the query, key, and value matrices, respectively, and $d_{k}$ is the dimension of the key vectors. This mechanism is applied within each domain’s sequence processing to capture both intra-domain and cross-domain relationships more effectively.
The framework generates nine sequence representations through multiple attention layers:
$h^{X}_{id},h^{X}_{img},h^{X}_{tex},h^{Y}_{id},h^{Y}_{img},h^{Y}_{tex},h^{X+Y}_{id},h^{X+Y}_{img},h^{X+Y}_{tex}.$ (5)
These representations capture both intra-domain and cross-domain user preferences. For domain $X$, the prediction probability for the next item $x_{t+1}$ in sequence $S^{X}$ is computed as follows:
$\mathrm{P}^{X}_{id}(x_{t+1}\mid\mathcal{S})=\text{softmax}\left(\text{Sim}(h_{id}^{X},E_{id}^{X})\right),\quad x_{t}\in S^{X},$ (6)
where $x_{t+1}$ denotes the predicted next item, $h_{id}^{X}$ is the ID-based sequence representation, and $E_{id}^{X}$ is the ID embedding matrix for domain $X$. Similarly, the prediction probabilities based on visual and textual features are calculated as:
$\mathrm{P}^{X}_{img}(x_{t+1}\mid\mathcal{S})$ $=\text{softmax}\left(\text{Sim}(h_{img}^{X},E_{img}^{X})\right),\quad$ $x_{t}\in S^{X},$ (7)
$\mathrm{P}^{X}_{tex}(x_{t+1}\mid\mathcal{S})$ $=\text{softmax}\left(\text{Sim}(h_{tex}^{X},E_{tex}^{X})\right),\quad$ $x_{t}\in S^{X},$
where $h_{img}^{X}$, $E_{img}^{X}$, $h_{tex}^{X}$, and $E_{tex}^{X}$ represent the image-based sequence representation, image embedding matrix, text-based sequence representation, and text embedding matrix, respectively. The final prediction probability for the next item combines ID, image, and text-based predictions:
$\mathrm{P}^{X}(x_{t+1}\mid\mathcal{S})$ $=\alpha\mathrm{P}^{X}_{id}(x_{t+1}\mid\mathcal{S})$
$\quad+\beta\mathrm{P}^{X}_{img}(x_{t+1}\mid\mathcal{S})$
$\quad+(1-\alpha-\beta)\mathrm{P}^{X}_{tex}(x_{t+1}\mid\mathcal{S}),$ (8)
where $\alpha$ and $\beta$ are weighting factors that balance the contributions of ID, visual, and textual features. This formulation captures structural, perceptual, and semantic aspects of user preferences.
To optimize the framework, we minimize the following loss function for domain $X$:
$\small\mathcal{L}^{X}=\sum_{x_{t}\in S^{X}}-\log\mathrm{P}^{X}(x_{t+1}\mid\mathcal{S}).$ (9)
The same method is applied to domains $Y$ and $X+Y$, yielding loss terms $\mathcal{L}^{Y}$ and $\mathcal{L}^{X+Y}$. The final loss is a weighted combination of these terms:
$\mathcal{L}=\mathcal{L}^{X}+\lambda_{1}\mathcal{L}^{Y}+\lambda_{2}\mathcal{L}^{X+Y},$ (10)
where $\lambda_{1}$ and $\lambda_{2}$ control the relative importance of each domain.
During evaluation, we aggregate predictions from all domains to recommend the next item. Let $X$ be the target domain and $Y$ the source domain. The combined prediction score for item $x_{i}$ is computed as:
$\mathrm{P}(x_{i}|S)=\mathcal{P}^{X}(x_{i}|S)+\lambda_{1}\mathcal{P}^{Y}(x_{i}|S)+\lambda_{2}\mathcal{P}^{X+Y}(x_{i}|S).$ (11)
The recommended item is selected by maximizing the prediction score within the target domain $X$:
$\mathrm{argmax}_{x_{i}\in\mathcal{X}}\mathrm{P}(x_{i}|S).$ (12)
This process allows the model to incorporate cross-domain information effectively, leveraging the predictions from both the target domain and the source domain. By weighting the contributions from each domain, the model can refine its recommendation, taking into account the richer information that might come from the source domain while focusing on the target domain for the final prediction.
The weighting factors $\lambda_{1}$ and $\lambda_{2}$ allow the model to control how much influence the source domain ($Y$) and the combined domain ($X+Y$) have on the recommendation for the target domain. This flexibility in adjusting the weights is crucial for handling different types of datasets or domain-specific characteristics, ensuring that the model adapts to varying degrees of relevance between domains.
Finally, by selecting the item with the highest prediction score from the target domain, the model aims to provide the most contextually appropriate recommendation, which is critical for real-world recommendation systems that need to balance both accuracy and relevance across multiple domains.
Table 1: Statistics of two CDSR scenarios.
Scenarios #Items #Train #Valid #Test Avg.length
Food 29,207 34,117 2,722 2,747 9.91
Kitchen 34,886 5,451 5,659
Movie 36,845 58,515 2,032 1,978 11.98
Book 63,937 5,612 5,730
Table 2: Experimental results (%) on the Food-Kitchen scenario.
Model (Food-Kitchen) MRR NDCG MRR NDCG
(Food) @5 @10 (Kitchen) @5 @10
NCF-MLP [71] 4.49 3.94 4.51 2.18 1.57 2.03
$\pi$-Net [5] 7.68 7.32 8.13 3.53 2.98 3.73
GRU4Rec [29] 5.79 5.48 6.13 3.06 2.55 3.10
SASRec [70] 7.30 6.90 7.79 3.79 3.35 3.93
SR-GNN [72] 7.84 7.58 8.35 4.01 3.47 4.13
PSJNet [73] 8.33 8.07 8.77 4.10 3.68 4.32
MIFN [74] 8.55 8.28 9.01 4.09 3.57 4.29
Tri-CDR [75] 8.35 8.18 8.89 4.29 3.63 4.33
MAN [14] 8.65 8.42 9.41 4.78 4.47 5.05
LLMRec [12] 9.05 8.86 10.03 5.02 4.73 5.34
IFCDSR [13] 9.05 8.85 9.92 4.95 4.56 5.24
Ours LLM-EMF 9.24 9.05 10.21 5.13 4.84 5.43
Table 3: Experimental results (%) on the Movie-Book scenario.
Model (Movie-Book) MRR NDCG MRR NDCG
(Movie) @5 @10 (Book) @5 @10
NCF-MLP [71] 3.05 2.26 2.96 1.43 1.06 1.26
$\pi$-Net [5] 4.16 3.72 4.17 2.17 1.84 2.03
GRU4Rec [29] 3.83 3.14 3.73 1.68 1.34 1.52
SASRec [70] 3.79 3.23 3.69 1.81 1.41 1.71
SR-GNN [72] 3.85 3.27 3.78 1.78 1.40 1.66
PSJNet [73] 4.63 4.06 4.76 2.44 2.07 2.35
MIFN [74] 5.05 4.21 5.20 2.51 2.12 2.31
MAN [14] 5.95 5.06 5.45 2.64 2.34 2.57
Tri-CDR [75] 5.15 4.62 5.05 2.32 2.08 2.22
LLMRec [12] 6.21 5.29 5.71 2.80 2.48 2.74
IFCDSR [13] 6.08 5.02 5.86 2.75 2.37 2.65
Ours LLM-EMF 6.32 5.35 5.78 2.86 2.53 2.79
4 Experiments
In this section, we introduce the experimental setup, providing detailed descriptions of the dataset, evaluation metrics, and implementation specifics. We then compare our method with state-of-the-art approaches on the Amazon dataset [76]. Finally, we conduct ablation studies to validate the effectiveness of the proposed method.
Dataset and Evaluated Metric. Our experiments are conducted on the Amazon dataset [76], which serves as the foundation for constructing cross-domain sequential recommendation (CDSR) scenarios. The dataset offers a rich collection of user-item interactions across a variety of domains, making it ideal for modeling cross-domain behaviors. Building on previous works [5, 74], we select four distinct domains to create two CDSR scenarios for our experiments: “Food-Kitchen” and “Movie-Book.” These domains were chosen to reflect real-world applications where users frequently engage with multiple categories, allowing us to investigate the efficacy of cross-domain knowledge transfer in sequential recommendation settings. The process of preparing the data begins by extracting users who have interacted with items in both domains. To ensure meaningful analysis, we filter out users and items with fewer than 10 interactions, focusing only on those with sufficient behavioral data. This filtering step is crucial for ensuring the robustness of our results, as sparse data can lead to unreliable recommendations. Furthermore, in order to meet the sequential constraints necessary for our study, we retain cross-domain interaction sequences that contain at least three items from each domain. This ensures that each sequence provides enough context for learning temporal dependencies and domain transitions. For the training/validation/test split, we adopt a temporal strategy to preserve the sequential nature of the data. Specifically, we divide the latest interaction sequences into validation and test sets, while the remaining sequences are allocated to the training set. This approach helps simulate a real-world scenario, where past user behavior informs future recommendations. The statistics of our refined datasets for the CDSR scenarios, including the number of users, items, and interactions in each domain, are summarized in Tab. 1.
To evaluate the performance of our model, we adopt two widely used evaluation metrics in recommendation systems: the Mean Reciprocal Rank (MRR) metric [77], which measures the rank of the first relevant item, and the Normalized Discounted Cumulative Gain (NDCG5, 10) metric  [78]. MRR is a metric that assesses the ranking quality by calculating the average of the reciprocal ranks of the first relevant item in each user’s recommendation list. NDCG, on the other hand, takes into account both the position and relevance of the recommended items, making it particularly suitable for evaluating ranked lists. These metrics are standard in the evaluation of recommendation systems, providing a comprehensive measure of recommendation quality. The results obtained from these metrics allow us to compare our model’s performance across different CDSR scenarios and demonstrate its effectiveness in capturing cross-domain interactions.
Figure 5: Impact of Hyperparameters on Final Performance on Movie Data
Implementation Details. For fair comparisons, we adopt the same hyperparameter settings as in previous works [74]. We conduct experiments with various hyperparameter settings, as shown in Figure 5. Specifically, we set the learnable embedding size $q$ to 256, the CLIP-based image and text embedding size $e$ to 512, and the mini-batch size to 256, with a dropout rate of 0.3. The values of $\lambda_{1}$ and $\lambda_{2}$ are set to 0.3 and 0.1, respectively. The $L_{2}$ regularization coefficient is selected from $\{0.00001,0.00005,0.0001,0.0005,0.001\}$, and the learning rate is chosen from $\{0.0001,0.0005,0.001,0.005,0.01\}$. Training is conducted for 100 epochs using an NVIDIA 4090 GPU, with the Adam optimizer [79] to update all parameters. To ensure robustness and reproducibility, we perform a grid search over the hyperparameter space, evaluating the model’s performance on a validation set at each step. The selected hyperparameters are those that yield the best validation performance. Additionally, we employ early stopping with a patience of 10 epochs to prevent overfitting, monitoring the validation loss during training. This approach ensures that the model generalizes well to unseen data while maintaining computational efficiency. The use of the Adam optimizer [79] is particularly advantageous due to its adaptive learning rate mechanism, which helps stabilize training and accelerate convergence. Furthermore, the NVIDIA 4090 GPU provides the necessary computational power to handle the large-scale datasets and complex model architectures efficiently. These settings collectively ensure that our experiments are conducted under optimal conditions, enabling meaningful comparisons with state-of-the-art methods in the field.
4.1 Performance Comparisons
We evaluate our proposed LLM-EMF method against a comprehensive set of state-of-the-art (SOTA) approaches on the “Food-Kitchen” and “Movie-Book” CDSR scenarios. The experimental results are summarized in Tab. 2 and Tab. 3, respectively. Our approach consistently outperforms all existing methods across both scenarios, demonstrating its effectiveness in handling cross-domain recommendation tasks. In the Food-Kitchen scenario, our LLM-EMF model achieves the highest scores across all evaluation metrics. Specifically, it attains an MRR of 9.24% for Food and 5.13% for Kitchen, significantly outperforming the closest competitor, MIFN, which achieves MRR scores of 8.55% and 4.09%, respectively. In terms of NDCG@5 and NDCG@10, LLM-EMF also leads with 9.05% and 10.21% for Food, and 4.84% and 5.43% for Kitchen. The performance of traditional methods such as ItemKNN, NCF-MLP, and BPRMF is notably lower, highlighting the limitations of these approaches in capturing complex user-item interactions in cross-domain settings. More advanced models like $\pi$-Net, GRU4Rec, and SASRec show improved performance but still fall short of LLM-EMF. This suggests that our model’s ability to leverage large language models (LLMs) for enhanced feature extraction and cross-domain knowledge transfer is a key factor in its superior performance. In the Movie-Book scenario, LLM-EMF again demonstrates its superiority, achieving MRR scores of 6.32% for Movie and 2.86% for Book. These results are significantly higher than those of the next best model, MIFN, which achieves MRR scores of 5.05% and 2.51%, respectively. Similarly, in terms of NDCG@5 and NDCG@10, LLM-EMF leads with 5.35% and 5.78% for Movie, and 2.53% and 2.79% for Book.
Table 4: Ablation study on the movie dataset
original-framework Textual Fusion LLM-enhanced Visual Fusion MRR 5.03 5.88 6.01 6.08 6.15 6.32
4.2 Ablation Studies
We conducted an ablation study to systematically evaluate the contributions of our proposed components: Textual Fusion, LLM-enhanced, and Visual Fusion. As illustrated in Tab. 4, we first established a baseline using the original framework without any additional modules, achieving an MRR of 5.03. The introduction of Textual Fusion alone resulted in a performance gain of 0.85 points, increasing the MRR to 5.88, which highlights the importance of leveraging textual information for cross-domain recommendation tasks. When the LLM-enhanced module was added alongside Textual Fusion, the MRR further improved to 6.01, demonstrating the effectiveness of leveraging large language models to enrich textual representations. Notably, the inclusion of Visual Fusion alone significantly boosted the MRR to 6.08, underscoring the critical role of visual signals in enhancing the model’s understanding of user preferences. The combination of Textual Fusion and Visual Fusion achieved an MRR of 6.15, indicating that textual and visual information complement each other in capturing cross-domain user behavior. Finally, the full framework, incorporating Textual Fusion, LLM-enhanced, and Visual Fusion, achieved the highest MRR of 6.32. This result demonstrates the cumulative impact of integrating multiple modalities and advanced language models for improving recommendation performance.
5 Conclusion
In this study, we propose a novel framework, named LLM-EMF, for cross-domain sequential recommendation. Unlike traditional methods that focus solely on temporal patterns within user interaction sequences, our approach leverages LLM knowledge to enrich textual information while also utilizing the powerful visual and textual understanding capabilities of the CLIP model to integrate multimodal representations. Specifically, we design prompts to combine the title and domain information, which are then used to generate augmented text. Both the images and augmented text are processed through CLIP to generate embeddings, which are subsequently fused with a learnable item matrix to enhance item representations. By combining these innovations, our LLM-EMF framework achieves state-of-the-art performance in CDSR, demonstrating the effectiveness of integrating multimodal signals and LLM knowledge in sequential recommendation tasks.
6 Acknowledgements
This work was supported by the National Natural Science Foundation of China (No. 62471405, 62331003, 62301451), Suzhou Basic Research Program (SYG202316) and XJTLU REF-22-01-010, XJTLU AI University Research Centre, Jiangsu Province Engineering Research Centre of Data Science and Cognitive Computation at XJTLU and SIP AI innovation platform (YZCXPT2022103).